OpenStudy (anonymous):

can some body help me please with alg 2?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i need help with that

OpenStudy (michele_laino):

hint: the first term is a_1= 3 -4 = -1 whereas the last term is a_18= 3*18 -4 = 50 there are 18 terms, so their sum is: \[{S_{18}} = \frac{{ - 1 + 50}}{2}18\]

OpenStudy (michele_laino):

\[{S_{18}} = \frac{{{a_1} + {a_{18}}}}{2}18 = \frac{{ - 1 + 50}}{2}18 = ...?\]

OpenStudy (anonymous):

umm not sure

OpenStudy (anonymous):

im not good at alg 2

OpenStudy (anonymous):

is it 49/262144